find the integral surface of the equation : (x-y)y^2p + (y- x)x^2q=(x^2 + y^2)z,
through the curve xz =a^2, y = 0
1
Expert's answer
2020-09-08T15:45:41-0400
"Solution"

The given partial differential equation is


"(x-y)y^2P+(y-x)x^2q=(x^2+y^2)z"


The Lagrange’s auxiliary equations for given PDE is

"\frac{dx}{(x-y)\cdot y^2}=\frac{dy}{(y-x)\cdot x^2}=\frac{dz}{(x^2+y^2) \cdot z}\\
Taking\ dx\ and\ dy\\
\frac{dx}{(x-y)\cdot y^2}=\frac{dy}{(y-x) \cdot x^2}"

"\frac{dy}{dx}=-\frac{x^2}{y^2}"

"dy \cdot y^2=dx \cdot (-x^2)"

"C_1=y^3+x^3------(i)"

"\frac{dx}{(x-y) \cdot y^2}=\frac{dz}{(x^2+y^2)\cdot z}"

"\frac{dz}{z}=\frac{dx(x^2+y^2)}{(x-y) \cdot y^2}"

"lnz=\frac{1}{y^2}\cdot F(x,y)+C_2--(ii)"


Here, y cannot be 0 in this equation that is why there is no intersection with curve "xz=a^2, y=0."


Answer: No solution.